OpenStudy (anonymous):

What is the simplified form of 4y^2-9 / 2y^2+y-3

OpenStudy (anonymous):



OpenStudy (anonymous):

4y^2-9 = (2y)^2 - 3^2 Use the identity \(a^2-b^2 = (a-b)(a+b)\) to factorize it. 2y^2 + y - 3 Split the middle term to make the factorization easier. = 2y^2 (- 2y + 3y) - 3 Factorize the first two and the last two terms respectively by taking out the common factors of terms. Can you do the above work first?

OpenStudy (anonymous):

Factorize the denominator, it's in the form ax^2 + bx + c

OpenStudy (anonymous):

can you show me how to factor the denominator, i just get stuck on that- I know the numerator

OpenStudy (anonymous):

So far, do you understand the step up to ''= 2y^2 (- 2y + 3y) - 3''?

OpenStudy (anonymous):

Yeah and i think i just did it ... would 2y+3 / y+1

OpenStudy (anonymous):

You can just use the cuadratic equation \[x = \frac{-b \pm \sqrt(b^2 -4ac)}{2a}\] I'd show you how in an easier way but I have to run to college

OpenStudy (anonymous):

with the restrction of y doesnt equal 3/2 and -1

OpenStudy (anonymous):

@kakrazz Not quite true, I mean the answer you get.

OpenStudy (anonymous):

ugh so the answer is wrong??

OpenStudy (anonymous):

What do you get for 2y^2 + y - 3 ? The factors, I mean.

OpenStudy (anonymous):

when i split the middle term, i do 2y^2 (-2y + 3y) - 3... but how would i factor that?

OpenStudy (anonymous):

how does that make it simpler?

OpenStudy (anonymous):

2y^2 + (-2y + 3y) - 3 ^ What are the common factors of the first two terms?

OpenStudy (anonymous):

the common factor of 2y^2 and -2y is 2...

OpenStudy (anonymous):

2y*

OpenStudy (anonymous):

Yes. So, take that out and group the rest for the first two terms, that is 2y^2 - 2y = 2y (y-1) Got it so far?

OpenStudy (anonymous):

yes but what do I do with the 3y and -3... OH WAIT DO I FACTOR BY GROUPING? omg im sorry! I know that now!

OpenStudy (anonymous):

i got it now! thank you @RolyPoly !!!!!

OpenStudy (anonymous):

Glad that you know it! What do you get for 2y^2 +y - 3 now?

OpenStudy (anonymous):

i get (2y+3)(y-1) and the (2y+3) will cancel out and the final denominator will be y-1 !!

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

THANK YOU!!!!

OpenStudy (anonymous):

Welcome! Make sure you know that NEITHER 2x+3 nor y-1 is equal to 0!

OpenStudy (anonymous):

YES! Thank you so so so so much!

OpenStudy (anonymous):

Welcome!